# Improved Contact Graph Routing in Delay Tolerant Networks with Capacity and Buffer Constraints

### v2

@@ II-D2 Bundle-dependent SR
III Contact-capacity and node-buffer challenges in DTNs
@@ IV Proposed global capacity and buffer management solution
Our solution builds upon bundle-dependent SR-CGR [31] (see Section II-D2). We address buffer and capacity constraints during the initial route-search phase. This proactive approach contrasts with existing reactive solutions that address issues after they occur.
Our approach guarantees that all discovered routes satisfy buffer and capacity constraints. The modifications also preserve all viable solutions, ensuring the discovery of optimal solution in the set of valid routes.
We achieve this by modifying the CP for both capacity and buffer constraints before route search begins. Additionally, we modify allowed edges between contacts during the route search to enforce buffer constraints. These modifications ensure CGR searches only within the space of valid solutions.
The system encodes the optimal route, which respects all buffer and capacity limits, into the bundle for network forwarding, following the standard source-routing CGR approach.
@@ IV-A Model-parameter modifications
  • First, we modify the CP through contact splitting for use in the route-search algorithm (i.e., Dijkstra search). The CP modifications fall into two categories:
@@
  • Second, we restrict edges between contacts through specific rules. The route-search algorithm enforces these rules during local route computation for each bundle.
@@ IV-A1 Contact splitting to modify the CP
We introduce the contact splitting operation. Consider contact $\text{C}_{j}$ with data rate $R_{\text{C}_{j}}$ between nodes $M$ and $N$, active from $t_{1,j}$ to $t_{2,j}$. Contact splitting divides this contact into two new contacts $\text{C}_{j,1}$ and $\text{C}_{j,2}$ between $M$ and $N$: one from $t_{1,j}$ to $t_{e_{1}}$ and another from $t_{e_{2}}$ to $t_{2,j}$, where $t_{1,j}\leq t_{e_{1}}\leq t_{e_{2}}\leq t_{2,j}$ (see Figure 3). Both split contacts maintain the original data rate $R_{\text{C}_{j}}$.
@@ IV-A2 Edge pruning
We define edge pruning as the act of forbidding edges between contacts. Edge pruning applies to any pair of contacts, whether original or split. In the example of Figure 2, one can choose to forbid the succesive use of contacts $\text{C}_{1}$ and $\text{C}_{2}$: pruning the edge between contacts $\text{C}_{1}$ and $\text{C}_{2}$ prevents their sequential use in any route.
Edge pruning occurs during the route search, specifically within the CRP. The implementation adds a new criterion to the contact exploration process (lines $8-12$ in Algorithm 1). When a contact fails this criterion, the algorithm:
@@ IV-B Capacity and buffer managements
We explain how contact splitting and edge pruning are used to address the capacity and buffer constraints discussed in Section III.
@@ IV-B1 Capacity management
This approach requires knowledge of each bundle’s size. The algorithm begins by searching a route for bundle $i$. Similarly to the linear-volume approach, the algorithm excludes contacts with insufficient volume for the bundle. Once the shortest path is selected for this bundle, the CP is modified as follows: We remove the part of the contact consumed by the bundle from the CP before performing the route search for subsequent bundles. The length of the contact to be erased is $t_{\text{Tx,}i}^{\text{C}_{j}}$, starting from the start time of the contact or the arrival time of the bundle at the contact’s sender node, whichever occurs later. The transmission time $t_{\text{Tx,}i}^{\text{C}_{j}}$ for bundle $i$ over contact Cj is:
$t_{\text{Tx,}i}^{\text{C}_{j}}=\frac{S_{i}}{R_{\text{C}_{j}}}.$ (1)
This contact splitting creates two new contacts (or one in case of shortening, see Section IV-A1) that replace the original one in the CP. We apply this operation to all contacts in the selected route. All subsequent route searches use this modified CP. These modifications are permanent and must propagate to all nodes performing route searches.
@@
Section VII shows our algorithm’s impact on the CP size, and thus on the CGR complexity, through practical examples.
@@ IV-B2 Buffer management
We introduce forecast buffer tables as a key component of our solution. These tables track buffer occupancy status for each node over time. Similarly to the CP, all route-computing entities share these tables. These tables then guide temporary modifications to both the CP and allowed edges.
Each node has a maximum buffer capacity $B_{\text{max}}$, with overflow occurring beyond this limit. Our buffer management solution applies temporary and bundle-specific CP modifications. Edge pruning follows the same temporary and bundle-specific pattern. These modifications remain local to the computing node and apply only during individual route searches.
Forecast buffer tables. Once a route is selected, the buffer tables are updated accordingly and shared to all entities performing a route search. In other words, once a route is validated, one checks when the bundle will reach a certain node and how long will it stay in its buffer according to the validated route.
@@
Algorithm to respect the buffer constrains. The proposed algorithm enforces buffer-limit constraints through the following approach: Using bundle size and forecast buffer tables, the algorithm verifies buffer availability at each node throughout the route. The algorithm excludes routes that would cause buffer overflow at any node. First, it modifies the CP temporarily through contact splitting. Next, it prunes specific edges during the route search. This algorithm can be formulated with the four following main steps.
@@
Example. Figure 4 presents an example where a bundle b arrives at a source node and needs to be routed. The figure depicts one node $Z_{1}\in\mathcal{Z}_{\text{b}}$ having two potential overflows $o_{1}\text{ and }o_{2}\in\mathcal{O}_{Z_{1}}$. These potential overflows are identified by virtually adding the size of bundle b (vertically striped rectangles) to the buffer level of node $Z_{1}$ (dotted rectangles). The buffer level of node $Z_{1}$, before virtually adding the size of bundle b, is determined by validated routes of previous bundles (dotted rectangles). The corresponding forecast buffer table for node $Z_{1}$ is shown in Figure 5.
@@
Step 4: Finally, the last step is to forbid edges between contacts yielding problematic contact successions. In other words, the edge with the contacts starting too late are pruned. This edge pruning is performed during the route search, in the CRP: lines $8-12$ in Algorithm 1. Note that additional inputs were required for the CRP to make our edge pruning possible: line $2$ in Algorithm 1. In addition to its original inputs, the CRP takes the bundle $b$ and the overflows start and end times $T_{1,s}$ and $T_{2,s}$ at the receiver of the selected contact $\text{C}_{s}$, respectively.
Example. In the example of Figure 6, during the route search, if a contact $\text{C}_{s}$ forwards the bundle to node $Z_{1}$ before $o_{1}$, the edge between $\text{C}_{s}$ and a contact $\text{C}_{i}$ from $Z_{1}$ to a third node is only accepted if $\text{C}_{i}$ starts before $t_{1,o_{1}}$. The same applies to all overflows at all nodes. More precisely, let us consider, in a CRP iteration, that $\text{C}_{s}$ is the selected contact and $\text{C}_{i}$ is another contact in the CP. Also, let us consider that the receiver of $\text{C}_{s}$ is node $Z_{1}$, having one potential overflow $o_{1}$. To consider $\text{C}_{i}$ as explored or a candidate successor for $\text{C}_{s}$, it must pass all the checks in the original CRP (lines $5-7$ in Algorithm 1). In addition to these checks, we add the following condition: If $t_{\text{arr},s}$ the bundle arrival time to the receiver of $\text{C}_{s}$ is before $t_{1,o_{1}}$ ($t_{\text{arr},s}<t_{1,o_{1}}$), then the start time of $\text{C}_{i}$ must also be before $t_{1,o_{1}}$ ($t_{1,i}<t_{1,o_{1}}$). Otherwise, $\text{C}_{i}$ is ignored in the CRP and it is not explored. If this condition is not met, $\text{C}_{i}$ cannot be a candidate successor for $\text{C}_{s}$. In a more general scenario, where multiple potential overflow periods exist at a node buffer, a bundle arriving at the node via contact $\text{C}_{s}$ between two overflow periods should depart from the node using contact $\text{C}_{i}$ before the next overflow period begins (lines $8-12$ in Algorithm 1).
As described by the above steps, only invalid routes are removed by the modifications of the proposed contact capacity and buffer management solutions. The proposed modifications before and during the route search do not omit any feasible solution. Hence, the subsequent use of CGR gives the optimal route in the subset of valid solutions.
@@
IV-C Booking resources for forwarding safety margin

### v3

@@ II-E2 Bundle-dependent SR
III Contact-capacity and node-buffer challenges in DTNs:
Presentation and problem formulation
@@ III-B Buffer management
III-C Mathematical problem formulation
We first formalize the constraints to be respected and then state the optimization problem. Given a path $P=(\text{C}_{1},\ldots,\text{C}_{k})$, with a transmission start time $\tau_{i}$ on contacts $\text{C}_{i}$, a given bundle size $S>0$, and release time $t_{0}$ (and source $s$, destination $d$):
III-C1 Contact-capacity feasibility
$P$ is contact-capacity feasible if it does not use any portion of a contact that is already allocated to another bundle (i.e., each contact used has a sufficient unallocated portion for the bundle).
III-C2 Buffer feasibility
Given $P$, the bundle is stored at node $v_{i}$ during the time interval
$\begin{split}&I_{v_{i}}=[\,\tau_{i}+S/R_{i}+\delta_{i},\ \tau_{i+1}\,]\ \ (i=1,\ldots,k-1).\end{split}$ (2)
Buffer feasibility requires (assuming no storage issue at the source node $s$ and at the destination node $d$)
$\begin{split}&S\leq\bar{B}_{v_{i}}(t)\quad\Leftrightarrow\quad S\leq\inf_{t\in I_{v_{i}}}\bar{B}_{v_{i}}(t),\\ &\forall t\in I_{v_{i}}\ (i=1,\ldots,k-1).\end{split}$ (3)
III-C3 Feasible route set
A feasible route is a path $P$ in the connectivity graph $(V_{c}^{0},E_{c}^{0})$ that is (i) schedulable, (ii) contact-capacity feasible, and (iii) buffer feasible. We denote the set of feasible routes by $\mathcal{P}_{f}(S,t_{0})$:
[…]
We denote the arrival time at destination as
$A(P)=\tau_{k}+S/R_{k}+\delta_{k}=f_{i_{k}}.$ (5)
With nonnegative traversal times ($S/R_{i}+\delta_{i}\geq 0$) and waiting allowed ($\tau_{i_{m+1}}\geq f_{i_{m}}$), the time-dependent graph is FIFO: on every (feasible) edge, a later departure cannot yield an earlier arrival at the receiver. We introduce this property so that the time-dependent shortest-path (TDSP) framework of [18] applies when establishing optimality (Section IV-C).
III-C4 Problem to solve: Feasible Earliest-Arrival Path with Capacity and Buffer constraints (FEAP-CB)
For a given bundle, the optimization problem is to find $P^{*}$ minimizing the arrival time among feasible routes $\mathcal{P}_{f}(S,t_{0})$:
$P^{*}=\operatorname*{arg\,min}\{\,A(P)\ :\ P\in\mathcal{P}_{f}(S,t_{0})\,\}.$ (6)
@@ IV Proposed global capacity and buffer management solution
Our solution builds upon bundle-dependent SR-CGR [6] (see Section II-E2). We address buffer and capacity constraints during the initial route-search phase. This proactive approach contrasts with existing reactive solutions that address issues after they occur.
Our approach guarantees that all discovered routes satisfy buffer and capacity constraints. The modifications also preserve all viable solutions, ensuring the discovery of an optimal solution in the set of valid routes (see Section IV-C for the proof).
The system then encodes the optimal route, which respects all buffer and capacity limits, into the bundle for network forwarding, following the standard source-routing CGR approach.
@@ IV-A Model-parameter modifications
  • We modify the CP through contact splitting for use in the route-search algorithm (i.e., Dijkstra search). The CP modifications fall into two categories:
@@
  • Second, we restrict feasible edges between contacts through specific rules. The route-search algorithm enforces these rules during local route computation for each bundle.
@@ IV-A1 Contact splitting to modify the CP
We introduce the contact splitting operation. Consider contact $\text{C}_{j}$ with data rate $R_{j}$ between nodes $M$ and $N$, active from $t_{1,j}$ to $t_{2,j}$. Contact splitting divides this contact into two new contacts $\text{C}_{j,1}$ and $\text{C}_{j,2}$ between $M$ and $N$: one from $t_{1,j}$ to $t_{e_{1}}$ and another from $t_{e_{2}}$ to $t_{2,j}$, where $t_{1,j}\leq t_{e_{1}}\leq t_{e_{2}}\leq t_{2,j}$ (see Figure 3). Both split contacts maintain the original data rate $R_{j}$.
@@
As a result, contact splitting (for capacity and for buffer) turns $V_{c}^{0}$ into the vertex set $V_{c}$ of split contacts.
@@ IV-A2 Edge pruning
We define edge pruning as the act of removing from the connectivity-only graph $(V_{c},E_{c}^{0})$ the edges that would violate buffer constraints. The remaining edge set is $E_{c}\subseteq E_{c}^{0}$. Edge pruning applies to any pair of contacts, whether original or split. For example, forbidding a given contact succession (e.g., the edge between $\text{C}_{1}$ and $\text{C}_{2}$) prevents their sequential use in any route.
Edge pruning occurs during the route search, specifically within the CRP. The implementation adds a new criterion to the contact exploration process (lines 8128-12 in Algorithm 1). When a contact fails this criterion, the algorithm:
@@ IV-B Capacity and buffer managements
We explain how contact splitting (for capacity and for buffer) and edge pruning (for buffer) are used to address the capacity and buffer constraints discussed in Section III.
The main idea is that a new contact graph $(V_{c},E_{c})$ is built by contact splitting and edge pruning. It will be shown (Lemma 3, Section IV-C) that the set of feasible routes $\mathcal{P}_{f}(S,t_{0})$ coincides with the set of schedulable paths in $(V_{c},E_{c})$. Running CGR on $(V_{c},E_{c})$ then provides the optimal path $P^{*}$.
@@ IV-B1 Capacity management
In terms of the contact graph, capacity-feasibility enforcement uses contact splitting to update the vertex set: from the current set of contact segments $V_{c}^{0}$ to a new set $V_{c}$ in which the portions already consumed by routed bundles are removed.
This approach requires knowledge of each bundle’s size $S$. The algorithm begins by searching a route for bundle $i$. Similarly to the linear-volume approach, the algorithm excludes contacts with insufficient volume for the bundle. Once the shortest path is selected for this bundle, the CP is modified as follows: We remove the part of the contact consumed by the bundle from the CP before performing the route search for subsequent bundles. In the contact-graph view, this contact splitting updates $V_{c}^{0}$ to $V_{c}$ (only contact segments with sufficient remaining capacity remain in $V_{c}$). The length of the contact to be erased is $t_{\text{Tx,}i}^{\text{C}_{j}}$, starting from the contact start time $t_{1,j}$ or the arrival time of the bundle at the contact’s sender node, whichever occurs later. The transmission time $t_{\text{Tx,}i}^{\text{C}_{j}}$ for bundle $i$ over contact $\text{C}_{j}$ is:
$t_{\text{Tx,}i}^{\text{C}_{j}}=\frac{S_{i}}{R_{j}}.$ (7)
This contact splitting creates two new contacts (or one in case of shortening, see Section IV-A1) that replace the original one in the CP. We apply this operation to all contacts in the selected route. All subsequent route searches use this modified CP (i.e., the updated vertex set $V_{c}$). These modifications are permanent and must propagate to all nodes performing route searches.
@@
Section VII shows the algorithm’s impact on the CP size, and thus on the CGR complexity, through practical examples.
@@ IV-B2 Buffer management
We assume that each node has a maximum buffer capacity $B_{\text{max}}$, with overflow occurring beyond this limit. Our buffer management solution applies temporary and bundle-specific CP modifications. Edge pruning follows the same temporary and bundle-specific pattern. These modifications remain local to the computing node and apply only during individual route searches.
Forecast buffer tables. We introduce forecast buffer tables as a key component of our solution. Once a route is selected, the buffer tables are updated accordingly and shared to all entities performing a route search. These tables record, for each node $n$ and time $t$, the residual buffer capacity $\bar{B}_{n}(t)$, or equivalently the booked load $Q_{n}(t)$. Similarly to the CP, all route-computing entities share these tables. They guide temporary modifications to both the CP and allowed feasible edges. In other words, once a route is validated, one checks when the bundle will reach a certain node and how long will it stay in its buffer according to the validated route.
@@
Algorithm to respect the buffer constraints. The proposed algorithm enforces buffer-limit constraints through the following approach: Using bundle size and forecast buffer tables, the algorithm verifies buffer availability at each node throughout the route. The algorithm excludes routes that would cause buffer overflow at any node. First, it modifies the CP temporarily through contact splitting (so $V_{c}^{0}$ becomes $V_{c}$). Next, it prunes specific feasible edges during the route search (removing from $E_{c}^{0}$ the edges that would cause overflow). This algorithm can be formulated with the four following main steps, see also Algorithm 2 for a summary. Then, the route search runs on $(V_{c},E_{c})$.
Algorithm 2 Buffer-constraint enforcement (summary)
1:Input: bundle $b$ (size $S$), forecast buffer tables, CP $(V_{c}^{0},E_{c}^{0})$
2:Output: feasible graph $(V_{c},E_{c})$ for route search
3:Step 1: Virtually add bundle size $S$ to all node buffer tables and detect potential overflows (nodes $\mathcal{Z}_{\text{b}}$, overflow sets $\mathcal{O}_{Z_{k}}$, times $t_{1,o_{l}}$, $t_{2,o_{l}}$).
4:Step 2: Temporarily modify CP: remove (via contact splitting) contact portions that would forward $b$ to any $Z_{k}\in\mathcal{Z}_{\text{b}}$ during an overflow interval and obtain vertex set $V_{c}$.
5:Step 3: Identify problematic contact successions (first contact delivers bundle to node before overflow, successor contact starts after overflow begins).
6:Step 4: Prune from $E_{c}^{0}$ the edges corresponding to those successions. Route search runs on $(V_{c},E_{c})$ (enforced in CRP, Algorithm 1, lines 8–12).
@@
Example.
Figure 4 presents an example where a bundle b arrives at a source node and needs to be routed. The figure depicts one node $Z_{1}\in\mathcal{Z}_{\text{b}}$ having two potential overflows $o_{1}\text{ and }o_{2}\in\mathcal{O}_{Z_{1}}$. These potential overflows are identified by virtually adding the size of bundle b (vertically striped rectangles) to the buffer level of node $Z_{1}$ (dotted rectangles). The buffer level of node $Z_{1}$, before virtually adding the size of bundle b, is determined by validated routes of previous bundles (dotted rectangles). The corresponding forecast buffer table for node $Z_{1}$ is shown in Figure 5.
@@
[…]
Example. In the example of Figure 6, during the route search, if a contact $\text{C}_{s}$ forwards the bundle to node $Z_{1}$ before $o_{1}$, the edge between $\text{C}_{s}$ and a contact $\text{C}_{i}$ from $Z_{1}$ to a third node is only accepted if $\text{C}_{i}$ starts before $t_{1,o_{1}}$. The same applies to all overflows at all nodes. More precisely, let us consider, in a CRP iteration, that $\text{C}_{s}$ is the selected contact and $\text{C}_{i}$ is another contact in the CP. Also, let us consider that the receiver of $\text{C}_{s}$ is node $Z_{1}$, having one potential overflow $o_{1}$. To consider $\text{C}_{i}$ as explored or a candidate successor for $\text{C}_{s}$, it must pass all the checks in the original CRP (lines 575-7 in Algorithm 1). In addition to these checks, we add the following condition: If $f_{s}$, the bundle arrival time to the receiver of $\text{C}_{s}$, is before $t_{1,o_{1}}$ ($f_{s}<t_{1,o_{1}}$), then the start time of $\text{C}_{i}$ must also be before $t_{1,o_{1}}$ ($t_{1,i}<t_{1,o_{1}}$). Otherwise, $\text{C}_{i}$ is ignored in the CRP and it is not explored. If this condition is not met, $\text{C}_{i}$ cannot be a candidate successor for $\text{C}_{s}$. In a more general scenario, where multiple potential overflow periods exist at a node buffer, a bundle arriving at the node via contact $\text{C}_{s}$ between two overflow periods should depart from the node using contact $\text{C}_{i}$ before the next overflow period begins (lines 8128-12 in Algorithm 1).
As described by the above steps, only invalid routes are removed and no feasible solution is omitted. Optimality of the route returned by CGR is formalized in Section IV-C.
@@
IV-C Optimality of the proposed approach
Lemma 1 (capacity enforcement). The capacity enforcement (contact splitting for capacity) modifies the contact graph so that a route is excluded (i.e., is no longer a path in the graph, because it would use a contact portion removed by splitting) iff it is capacity-infeasible (i.e., does not respect the contact-capacity constraint defining $\mathcal{P}_{f}$): (i) every route excluded in this way would violate that constraint, and (ii) every route that would violate it is excluded.
Proof. (i) If a route is excluded by the capacity enforcement, it would use at least one contact portion that was removed by contact splitting for capacity (already allocated to prior bundles). By construction, any such route would exceed available contact capacity. So every route excluded in this way is capacity-infeasible (does not respect the contact-capacity constraint defining $\mathcal{P}_{f}$). (ii) If a route would violate the contact-capacity constraint, then it would use a portion of some contact that is already allocated to another bundle. That contact portion is removed by contact splitting for capacity. Thus every capacity-infeasible route is excluded (no longer a path in the graph). $\square$
Lemma 2 (buffer enforcement, Algorithm 2). The buffer enforcement builds the feasible graph $(V_{c},E_{c})$: contact splitting for buffer updates the vertex set $V_{c}$, and edge pruning yields $E_{c}\subseteq E_{c}^{0}$. A route is excluded (i.e., is not a path in $(V_{c},E_{c})$) iff it is buffer-infeasible (i.e., does not respect the node-buffer constraint defining $\mathcal{P}_{f}$): (i) every route excluded in this way (it uses a contact portion removed by buffer splitting or an edge pruned for buffer) would violate that constraint, and (ii) every route that would violate it is excluded.
Proof. (i) The buffer procedure builds $(V_{c},E_{c})$ via contact splitting (Step 2) and edge pruning (Step 4). A route is excluded (not a path in $(V_{c},E_{c})$) only if it would use a contact portion forwarding $b$ to some $Z_{k}\in\mathcal{Z}_{\text{b}}$ during an overflow interval, or an edge corresponding to a problematic succession. By construction, any such route would exceed $\bar{B}$ at some node and time. So every route excluded in this way is buffer-infeasible (does not respect the node-buffer constraint defining $\mathcal{P}_{f}$). (ii) If a route would violate the node-buffer constraint at some node $Z$, then $Z\in\mathcal{Z}_{\text{b}}$ and the route either forwards $b$ to $Z$ during an overflow interval (hence that contact portion is removed in Step 2) or uses a problematic succession at $Z$ (hence the corresponding edge is pruned in Step 4). Thus every buffer-infeasible route is excluded (not a path in $(V_{c},E_{c})$). $\square$
Lemma 3 (bijection). A route is a feasible route (i.e., $P\in\mathcal{P}_{f}(S,t_{0})$ as defined in Section III-C) iff it is a schedulable vertex path in $(V_{c},E_{c})$.
Proof. ($\Rightarrow$) A feasible route is, by definition, in $\mathcal{P}_{f}(S,t_{0})$, hence capacity-feasible and buffer-feasible. By Lemmas 1 and 2, a route is excluded (no longer a path in the graph / not a path in $(V_{c},E_{c})$) iff it violates the contact-capacity or node-buffer constraint defining $\mathcal{P}_{f}$. So a feasible route is not excluded and uses only contact portions and successions that remain in $(V_{c},E_{c})$. Its vertices therefore lie in $V_{c}$ and its edges in $E_{c}$, and with its per-contact start times it is schedulable. Thus the route induces a schedulable vertex path in $(V_{c},E_{c})$. ($\Leftarrow$) A schedulable vertex path in $(V_{c},E_{c})$ uses only vertices in $V_{c}$ and edges in $E_{c}$, so it was not excluded by the capacity or buffer enforcement. By Lemmas 1 and 2, it is therefore capacity-feasible and buffer-feasible. With schedulability it lies in $\mathcal{P}_{f}(S,t_{0})$, i.e., it is a feasible route. $\square$
In the framework of [18], the following holds.
Lemma 4 (FIFO time-dependent shortest path). On a FIFO, nonnegative time-dependent graph with waiting allowed, time-dependent Dijkstra returns a path of minimum arrival time for a fixed departure time.
Theorem (optimality). Let $P^{*}$ be the vertex path returned by time-dependent Dijkstra on $(V_{c},E_{c})$. Then $A(P^{*})=\min\{\,A(P):P\in\mathcal{P}_{f}(S,t_{0})\,\}$.
Proof. By Lemma 4, $P^{*}$ minimizes arrival time among all vertex paths in $(V_{c},E_{c})$. By Lemma 3, those vertex paths are exactly the feasible routes (elements of $\mathcal{P}_{f}(S,t_{0})$). Hence $P^{*}$ is optimal. $\square$
IV-D Booking resources for forwarding safety margin